OpenStudy (anonymous):

Choose all properties that were used to simplify the following problem: 1. 2/7• 53 •7/2 2. 53 •2/7 •7/2 3. 53 • 1 4. 53 A. commutative property of addition B. commutative property of multiplication C. associative property of addition D. associative property of multiplication E. multiplicative identity F. multiplicative inverse G. distributive property

OpenStudy (anonymous):

ohh wait i think i have the notes...gimme a sec.

OpenStudy (anonymous):

okay man

OpenStudy (anonymous):

-Commutative property of multiplication ex: a*b=b*a -Associative property of multiplication ex: (n*n)+n or n+(n*n) -Distributive property ex: a(b+c)=a*b+a*c

OpenStudy (anonymous):

well right now you can just eliminate the ones that say addition bcz theres no need to add right now

OpenStudy (anonymous):

okay bro